Complementary Probabilities in Trials

00:00 2 marks

  • Build Confidence
  • SL
  • Tier 1
  • Probability
  • Complement rule

A certain event A has a probability of 0.25 of occurring in a single trial. Using the complement rule, find the probability that event A does not occur.

Hint Hint

Hint

Recall the rule: P(A’) = 1 – P(A).

  • Answer Answer
  • Mark Scheme Mark Scheme
  • Explanation Explanation
  • Common Errors Common Errors
  • Real World Context Real World Context
  • Focus Focus
  • Formula Booklet
  • Answer Answer

Complete Answer

0.75

  • Mark Scheme Mark Scheme
  • Answer Explanation Answer Explanation
  • Common Errors Common Errors
  • Focus Focus

Math Editor

Report Question


Please enter keywords